# Evolution of boundedly rational learning in games

### v1

@@ Abstract
People make strategic decisions multiple times a day. We act strategically in negotiations, when we coordinate our actions with others, or when we choose with whom to cooperate. The resulting dynamics can be studied with evolutionary game theory. This framework explores how people adapt their decisions over time, in light of how effective their strategies have proven to be. A crucial quantity in respective models is the strength of selection. This quantity regulates how likely individuals switch to a better strategy when given the choice. The larger the selection strength, the more biased is the learning process in favor of strategies with large payoffs. Therefore, this quantity is often interpreted as a measure of rationality. Traditionally, most models take selection strength to be a fixed parameter. Instead, here we allow the individuals’ strategies and their selection strength to co-evolve. The endpoints of this co-evolutionary process depend on the strategic interaction in place. In many prisoner’s dilemmas, selection strength increases indefinitely, as one may expect. However, in snowdrift or stag-hunt games, it can either converge to a finite value, or we observe evolutionary branching altogether – such that different individuals arrive at different selection strengths. Overall, this work sheds light on how evolution might shape learning mechanisms for social behavior. It suggests that boundedly rational learning is not only a by-product of cognitive constraints. Instead it might also evolve as a means to gain strategic advantages.
Keywords: evolutionary game theory |||| adaptive dynamics |||| learning |||| social dilemma
@@ Introduction
People constantly make strategic decisions. These decisions range from rather inconsequential choices (e.g., whether to do a small favor) to major commitments (e.g., choosing a business partner). The process by which individuals arrive at their decisions can be analyzed with evolutionary game theory [1, 2, 3]. This framework treats individuals as players who interact in games. They can choose among different strategies and win a payoff which depends both on their own decisions and those of their co-players. Importantly, evolutionary game theory does not require individuals to choose optimal decisions from the outset. Rather, individuals are assumed to engage in the same interaction (or sufficiently similar ones) many times. Over the course of these interactions, they adapt their behavior to yield higher payoffs. The respective dynamical process allows for several interpretations. We can think of strategies as being genetically determined and evolving across generations [4, 5]. Alternatively, models of learning consider how individuals switch between strategies within their lifetime [6, 7, 8, 9, 10, 11, 12, 13]. Typically, these two interpretations are studied separately or interchangeably. In practice, however, the two approaches are intertwined: social behavior is often shaped by learning, while learning mechanisms are shaped by evolutionary forces [14, 15, 16, 17].
@@
Herein, we ask how evolutionary forces might shape a key parameter of many stochastic learning models, the selection strength or intensity of selection. This parameter $\beta\!\geq\!0$ determines to which extent an individual’s choice to switch strategies is driven by payoff differences  (Fig. 1A). When this parameter is small, players are largely insensitive to their strategies’ payoffs. They mostly switch from one strategy to another by chance, with payoff differences only slightly affecting that probability (Fig. 1B, left). The respective limit of weak selection ($\beta\!\rightarrow\!0$) has become a major paradigm in evolutionary game theory, partly because it simplifies many calculations [18, 19, 20, 21, 22]. In this limit, learning takes the form of an (almost unbiased) random walk on the space of all strategies. In contrast, as selection strength becomes large, players only switch strategies if the new strategy is superior (Fig. 1B, right). The respective limit of strong selection ($\beta\!\rightarrow\!\infty$) has become another major paradigm to study evolutionary games [23, 24, 25]. As one advantage of that limit, respective results are closely aligned with key concepts in classical game theory. For example, in this limit strict Nash equilibria correspond to absorbing states of the learning process [26, 27]. Herein, we thus interpret the value of $\beta$ as a measure of how rational individuals are when updating their strategies [similar to the interpretation of a related quantity in the ‘logit equilibrium’ 28, 29, 30]. Accordingly, we think of finite $\beta$ values as indicating boundedly rational learning.
Figure 1: An overview of the model. A, In stochastic evolutionary game theory, individuals continually get a chance to revise their strategy. We model this revision process with introspection dynamics [31, 32, 33, 34, 35, 36, 37]. Here, a player compares its current strategy to a randomly chosen alternative. If $\pi$ is the payoff of the current strategy, and $\pi^{\prime}$ the payoff of the alternative, the player’s switching probability is given by a Fermi function $\phi(\pi^{\prime}\!-\!\pi)$. This function captures that players are more likely to switch if the alternative strategy is more beneficial. B, The exact shape of the switching probability depends on a parameter $\beta$, the strength of selection. For low $\beta$, payoff differences between strategies become irrelevant. Here, strategy changes are mostly driven by chance. As $\beta$ becomes larger, individuals increasingly favor those alternative strategies with high payoffs. Thus, $\beta$ is often interpreted as a measure of the players’ rationality or their learning effectiveness. We are interested in the evolution of this parameter. Our process unfolds on two timescales. C, In the short run, each player’s selection strength $\beta$ is fixed. Given their selection strength, they choose between different strategies in a stage game. The depicted example shows a round in which player 1 switches from strategy C (cooperation) to strategy D (defection). We iterate this learning process for many rounds. Based on these iterations, we compute the average frequency of each possible game outcome (illustrated as a black-and-white gradient). This allows us to compute the players’ expected payoff as a function of their $\beta$ values. D, In the long-run, we let the players’ selection strength $\beta$ evolve. We model this long-run process with adaptive dynamics [38, 39, 40, 41].
@@
Except for a handful of studies discussed further below [42, 43], most models interpret selection strength as a fixed quantity, which is the same for everyone. Here, we relax this assumption. We address two related questions. First, we are interested in learning processes among individuals with fixed but different selection strengths. We refer to individuals with larger $\beta$ as more effective learners, and those with lower $\beta$ as less effective learners. We ask to which extent more effective learners are able to systematically outperform their opponents, by choosing their strategies more optimally.
@@
Second, we explore the dynamics when an individual’s selection strength is itself an evolving trait. To this end, we use a framework akin to the indirect evolutionary approach used to study the evolution of preference [44, 45, 46, 47, 48]. The dynamics unfolds on two time scales. In the short run, the players’ selection strengths are fixed. Players interact in a given game with others, adopt new strategies based on their selection strength, and as a result, earn a certain average payoff (Fig. 1C). In the long run, we allow the selection strength to evolve, based on the average payoffs the respective players yield. To the extent that selection strength is an inheritable trait, this assumption may reflect that more effective learners have more offspring (Fig. 1D). We explore this long-run dynamics analytically and with simulations, based on the adaptive dynamics formalism [38, 39, 40, 41]. To get an intuitive understanding, we focus in the main text on the most simple class of games: symmetric games among two players with two strategies. These games include, for example, the prisoner’s dilemma, the stag-hunt game, or the snowdrift game. However, we also show that the respective intuitions carry over to more general multiplayer games.
@@
Once a player’s selection strength is an evolvable trait, one might expect an evolutionary dynamics towards ever-increasing selection strengths. After all, players with larger selection strength are better at identifying strategies with high payoffs. Hence, such players should have a fitness advantage. Remarkably, however we do not find such a runaway dynamics for all games. Instead in many snowdrift games, selection strength converges to a finite value. Moreover, in many stag-hunt games we observe evolutionary branching. Here, an initially monomorphic population splits into two or more subpopulations, each with a different selection strength. Interestingly, we obtain these results even without assuming that more effective learners need to exert more cognitive effort and hence pay some complexity cost. Instead, we obtain these results because an ineffective learning process can turn out to be a strategic advantage. In this way, our study contributes to a growing literature exploring how strategic considerations can explain behaviors that are seemingly at odds with individually rational decision-making [44, 45, 46, 47, 48].
@@
Our model assumes a separation of timescales. In the short run, players have fixed but possibly different selection strengths. They repeatedly engage in a given one-shot game, and they learn to adopt more profitable strategies over time. We model this short-run learning process with introspection dynamics, a process particularly useful to study games with some asymmetry [31, 32, 33, 34, 35, 36, 37]. In the long run, the players’ selection strengths are allowed to evolve. We describe this long-run process with adaptive dynamics [38, 39, 40, 41], a classical tool to study the evolution of continuous traits. In the following, we describe the two processes in turn.
@@
To explore learning among players with different selection intensities, we consider simple $2\times 2$ games. That is, we consider two players who can choose among two possible strategies, $\mathbf{C}$ and $\mathbf{D}$ (we think of the two strategies as cooperation and defection, but this interpretation is irrelevant for our results). The choices of the two players determine their payoffs. Payoffs are assumed to be symmetric, and given by the (normalized) payoff matrix
$\begin{array}[]{c|cc}&\mathbf{\;C\;}&\mathbf{\;D\;}\\ \hline\cr\mathbf{\;C\;}&1&-B\\ \mathbf{\;D\;}&1-A&0\\ \end{array}$ (1)
We refer to this matrix as the stage game. As usual, player 1 chooses a row of this matrix, player 2 chooses a column, and the entries represent the payoffs of player 1. Depending on the signs of $A$ and $B$, we recover four major game classes: prisoner’s dilemma ($A\!<\!0$, $B\!>\!0$), stag-hunt game ($A\!>\!0$, $B\!>\!0$), snowdrift game ($A\!<\!0$, $B\!<\!0$), and harmony game ($A\!>\!0$, $B\!<\!0)$.
To describe how players learn to adopt more profitable strategies over time, we use introspection dynamics [31, 32]. This process assumes that the two individuals interact in the same stage game over many time steps. Each time step, a randomly chosen player is given a chance to revise its strategy. That player then compares its current payoff $\pi$ with the payoff $\pi^{\prime}$ the player could have obtained by choosing the other strategy (keeping the co-player’s strategy fixed). As illustrated in Fig. 1A, the player switches to the other strategy with a probability that depends on the payoff difference [6, 7, 8],
@@
The parameter $\beta_{i}\!\geq\!0$ represents player $i$’s selection strength. For small $\beta_{i}$, this function approaches $1/2$ irrespective of the payoff difference (Fig. 1B left). In this limit, players decide whether to switch strategies based on a coin toss. For large $\beta_{i}$, the function approaches a step function (Fig. 1B right). Here, the player always switches if the alternative strategy yields the better payoff. We iterate this basic updating process for many time steps. As a result, players repeatedly get a chance to switch their strategies  (Fig. 1C, bottom). Based on this process, we record how frequently we observe each of the four possible game outcomes over time. In the right panel of Fig. 1C, we illustrate these overall frequencies using a black-and-white gradient. A darker shading indicates game outcome that are observed more often.
Introspection dynamics is particularly convenient because the expected frequencies of the four game outcomes can be computed analytically. To this end, we represent the learning dynamics as a Markov chain (see Supporting Information for details). The Markov chain has four states, corresponding to the four possible outcomes in each time step, CC, CD, DC, DD (the first letter refers to the first player’s strategy, and the second letter to the second’s). For finite $\beta_{1},\beta_{2}$, this Markov chain has a unique stationary distribution $\mathbf{u}(A,B,\beta_{1},\beta_{2})\leavevmode\nobreak\ :=\leavevmode\nobreak\ (u_{\textbf{CC}},u_{\textbf{CD}},u_{\textbf{DC}},u_{\textbf{DD}})$. This distribution captures how often we observe each of the four game outcomes on average, over the course of many rounds. Once we have the stationary distribution, it becomes straightforward to compute player 1’s average payoff, by multiplying the four frequencies $u_{ij}$ with the respective stage game payoffs,
@@
The respective results become particularly simple when both players have the same selection strength, $\beta_{1}\!=\!\beta_{2}$. In that case, one can show that the stationary distribution simplifies to
@@
For weak selection ($\beta\!\rightarrow\!0$), this formula implies that all game outcomes are observed equally often, $\mathbf{u}=(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})$. For strong selection ($\beta\!\rightarrow\!\infty$), the stationary distribution depends on the game parameters $A$ and $B$. For example, for the prisoner’s dilemma, (4) implies that the two players defect in almost all rounds ($u_{\textbf{DD}}\!=\!1$); in the snowdrift game, they anti-coordinate ($u_{\textbf{CD}}\!=\!u_{\textbf{DC}}\!=\!\frac{1}{2}$); and in the stag-hunt game they choose the strategy that is risk-dominant [$u_{\textbf{DD}}\!=\!1$ if $A\!<\!B$, $u_{\textbf{CC}}\!=\!1$ if $A>B$, see Ref. 31].
Using the same formalism, we can also explore the case when players have different selection strengths. In the Supporting Information, we derive a general expression of the stationary distribution $\mathbf{u}(A,B,\beta_{1},\beta_{2})$ as a function of $\beta_{1}$ and $\beta_{2}$, which is now more complex. In the following, we use that expression to explore whether a more effective learner (one with larger $\beta$) would gain a payoff advantage in various stage games.
In Fig. 2A–D, we consider a prisoner’s dilemma. We assume the two players can either have a small or a comparably large selection strength, $\beta\!\in\!\{2,5\}$. For a symmetric situation in which both players use the same $\beta$, we find that players learn to defect more often when $\beta$ is large (Fig. 2A,B). This result is intuitive: the larger the selection strength, the better players are at identifying the individually optimal strategy of defecting. Once players are asymmetric, it is the more effective learner who defects more often (Fig. 2C). As a result, this player ends up with a better average payoff.
@@
Figure 2: Introspection dynamics among players with different selection strengths. To illustrate the impact of selection strength on the learning dynamics, we consider two stage games, a prisoner’s dilemma (A–C) and a snowdrift game (D–F). In each case, we assume selection strengths may assume two possible values. They can either be comparably low (bright avatar) or high (dark avatar). For all possible combinations of selection strengths, we depict how often the respective players would obtain one of the four possible payoffs of the respective stage game. Based on this stationary distribution, we compute the players’ expected payoffs. For the prisoner’s dilemma we observe that the higher selection strength dominates the lower selection strength (panel D). In contrast, for the snowdrift game, it is the low-selection strength player who gains the larger average payoff (panel H).
In Fig. 2E–H, we consider a snowdrift game. For illustration, the possible selection strengths are now $\beta\!\in\!\{5,15\}$. In the symmetric case among players with equal selection strength, a large $\beta$ again makes it more likely that players implement a Nash equilibrium (here, CD or DC). In the depicted example, they coordinate on a Nash equilibrium in $74\%$ of all rounds when $\beta\!=\!5$, but in $90\%$ of all rounds when $\beta\!=\!15$. Remarkably, however, once players have different selection strengths, it is the less effective learner who obtains the larger average payoff. By acting myopically against a comparably rational opponent, such players ensure that their more preferred equilibrium is visited more often (CD in Fig. 2G). In this way, they obtain a superior payoff – $\Pi_{15}(5)\!=\!0.86>0.52\!=\!\Pi_{5}(15)$ – despite being less sophisticated in their strategy choices.
@@
Based on these results, we can construct a ‘super-game’ matrix (Fig. 2D,H). This matrix represents the interaction of two learners who choose between different selection strengths. The entries of the matrix depict the learners’ average payoffs over the course of the learning process (as derived in the previous panels). For the prisoner’s dilemma, the super-game indicates that each learner always prefers to have a higher selection strength. In the snowdrift game example, we observe the opposite. Here, it is the lower selection strength that dominates the higher one.
@@
Overall, these two examples provide a proof of principle. They illustrate the non-trivial effects of asymmetric selection strengths. These observations call for a more systematic analysis, considering more games and a larger set of possible selection strengths. We provide such an analysis in Fig. S1 of the Supporting Information. In addition, the above results raise the question whether selection strength would necessarily move towards ever-increasing values if it was an evolving trait. We explore that issue in the next section.

### v2

# Evolution of noisy learning in games

@@ Abstract
People make strategic decisions many times a day – during negotiations, when coordinating actions with others, or when choosing partners for cooperation. The resulting dynamics can be studied with learning theory and evolutionary game theory. These frameworks explore how people adapt their decisions over time, in light of how effective their strategies have been. The outcomes of such learning processes depend on how sensitive individuals are to the performance of their strategies. When they are more sensitive, they systematically favor strategies they deem more successful. When they are less sensitive, their learning process is noisier and more erratic. Traditionally, most models treat this sensitivity as a fixed parameter – like the ‘selection strength’ parameter in evolutionary models. Instead, we study how strategies and sensitivities co-evolve. We find that the co-evolutionary endpoints depend on both the type of strategic interaction and the learning rule employed. In prisoner’s dilemmas, we often observe sensitivities to increase indefinitely. But in snowdrift and stag-hunt games, sensitivities often converge to a finite value, or we observe evolutionary branching altogether. These results shed light on how evolution might shape learning mechanisms for social behavior. They suggest that noisy learning does not need to be a by-product of cognitive constraints. Instead, it can serve as a means to gain strategic advantages.
Keywords: evolutionary game theory $|$ adaptive dynamics $|$ learning $|$ social dilemma
@@ Introduction
People constantly make strategic decisions. These decisions range from inconsequential choices (e.g., whether to do a small favor) to major commitments (e.g., choosing a business partner). The process by which individuals arrive at their decisions can be analyzed with learning theory [38, 23, 21] or evolutionary game theory [45, 12, 60, 43, 42, 37]. To this end, individuals are regarded as players who interact in games. They can choose among different strategies and win a payoff that depends both on their own and their co-players’ decisions. Importantly, they do not need to behave optimally from the outset. Rather, individuals are assumed to engage in the same interaction (or sufficiently similar ones) many times. Over the course of these interactions, they can adapt their behavior to yield better payoffs.
@@
The outcome of this adaptation process depends on how sensitive players are to their strategies’ performance. Most models capture this sensitivity by a parameter. For example, in stochastic learning models of evolutionary game theory, outcomes crucially depend on the selection strength or intensity of selection, $\beta$. This parameter $\beta$ determines to which extent individuals switch their strategies because of payoff differences (Fig. 1A). When this parameter is small, players are largely insensitive to their payoffs. They mostly switch from one strategy to another by chance (Fig. 1B, left). The respective limit of weak selection ($\beta\!\rightarrow\!0$) has become a major paradigm in evolutionary game theory, partly because it simplifies many calculations [44, 58, 41, 3, 54]. In this limit, learning takes the form of an (almost unbiased) random walk on the space of all strategies. In contrast, as selection strength becomes large, players only switch strategies if the new strategy is superior (Fig. 1B, right). The respective limit of strong selection ($\beta\!\rightarrow\!\infty$) has become another major paradigm to study evolutionary games [30, 53, 24]. Respective results are closely aligned with key concepts in classical game theory. For example, in this limit, strict Nash equilibria correspond to absorbing states of the learning process [22, 20]. In the following, we refer to $\beta$ as a player’s (payoff) sensitivity. A small $\beta$ translates into a more noisy learning process. A large $\beta$ makes individuals update their strategies more deterministically. Related parameters also exist in other learning models. For example, in experience-weighted attraction learning [13, 19, 48], outcomes critically depend on a player’s sensitivity to attractions.
@@
Except for a handful of studies discussed further below [56, 57], most models interpret a player’s (payoff) sensitivity as a fixed quantity, often assumed to be the same for everyone. Here, we relax this assumption. We address two related questions. First, we are interested in learning processes among individuals with fixed but different sensitivities. We refer to individuals with larger $\beta$ as more sensitive learners, and those with lower $\beta$ as less sensitive learners. We ask to what extent more sensitive learners are able to systematically outperform their opponents.
Figure 1: An overview of the model. A, In stochastic models of evolutionary game theory, individuals continually get a chance to revise their strategy. We model this revision process with introspection dynamics [15, 16, 32, 40, 63, 36, 52, 50]. Here, a player compares its current strategy to a randomly chosen alternative. If $\pi$ is the payoff of the current strategy, and $\pi^{\prime}$ the payoff of the alternative, the player’s switching probability is given by a Fermi function $\phi(\pi^{\prime}\!-\!\pi)$. This implies players are more likely to switch if the alternative strategy is more beneficial. B, The exact shape of the switching probability depends on a parameter $\beta$. This parameter is often referred to as the strength of selection. Because here we interpret $\beta$ as an individual (and evolvable) trait, we refer to it as the player’s payoff sensitivity. For low $\beta$, the player’s learning process is more noisy. Here, strategy changes are mostly driven by chance. As $\beta$ becomes larger, updating decisions become more deterministic. Here, individuals increasingly favor those alternative strategies with high payoffs. We are interested in the evolution of this parameter $\beta$. Our process unfolds on two timescales. C, In the short run, each player’s payoff sensitivity $\beta$ is fixed. Given their sensitivity, they choose between different strategies in a stage game. The depicted example shows a round in which player 1 switches from strategy C (cooperation) to strategy D (defection). We iterate this learning process for many rounds. Based on these iterations, we compute the average frequency of each possible game outcome (illustrated as a black-and-white gradient). This allows us to compute the players’ expected payoff as a function of their $\beta$ values. D, In the long-run, we let the players’ payoff sensitivity $\beta$ evolve. We model this long-run process with adaptive dynamics [27, 26, 34, 11].
@@
Second, we explore the dynamics when a player’s sensitivity is itself an evolving trait. To this end, we consider a framework akin to the indirect evolutionary approach used to study the evolution of preference [28, 35, 1, 2, 62]. The dynamics unfolds on two time scales. In the short-run, sensitivities are fixed. Players interact in a given game with others, adopt new strategies based on their sensitivity, and as a result, earn a certain average payoff (Fig. 1C). In the long-run, we allow their sensitivity to evolve, based on the average payoffs they yield. To the extent that a player’s sensitivity is an inheritable trait, this assumption may reflect that more effective learners have more offspring (Fig. 1D). We explore this long-run dynamics analytically and with simulations, based on the adaptive dynamics formalism [27, 26, 34, 11].
To get an intuitive understanding, we focus on the simplest class of games: symmetric games among two players with two strategies. These games include the prisoner’s dilemma, the stag-hunt game, and the snowdrift game. However, we also show that the respective intuitions carry over to more general settings, including multiplayer and repeated games. Similarly, throughout most of the main text, we study a baseline model based on one particular learning rule, called introspection dynamics [15, 16, 32, 40, 63, 36, 52, 50]. This dynamics is particularly convenient to work with because it allows for a fully analytical treatment. However, further below (and in more detail in the Supporting Information), we discuss to which extent the respective results generalize to other learning rules.
@@
Once a player’s sensitivity is an evolvable trait, one might expect an evolutionary dynamics towards ever-increasing sensitivities. After all, more sensitive learners are better at identifying strategies with high payoffs. Hence, such learners might have a fitness advantage. Remarkably, however, we do not find such a runaway dynamics for all games. Instead, in many snowdrift games, the players’ sensitivities converge to a finite value. Moreover, in some stag-hunt games, we observe evolutionary branching. Here, an initially monomorphic population splits into two or more subpopulations, each with a different sensitivity. Interestingly, we obtain these results without assuming that more sensitive learners exert more cognitive effort and hence pay some complexity cost. Instead, we obtain these results because a noisy learning process can turn out to be a strategic advantage. In this way, our study contributes to a growing literature exploring how strategic considerations can explain behaviors that are seemingly at odds with sensible decision-making [28, 35, 1, 2, 62].
@@
Our model assumes a separation of timescales. In the short run, players have fixed but possibly different payoff sensitivities. They repeatedly engage in a given one-shot game, and they learn to adopt more profitable strategies over time. We model this short-run learning process with introspection dynamics [15, 16, 32, 40, 63, 36, 52, 50]. In the long run, the players’ payoff sensitivities are allowed to evolve. We describe this long-run process with adaptive dynamics [27, 26, 34, 11], a classical tool to study the evolution of continuous traits. In the following, we describe the two processes in turn.
@@
To explore learning among players with different payoff sensitivities, we consider simple one-shot, $2\times 2$ games. That is, we consider two players who can choose among two possible strategies, $\mathbf{C}$ and $\mathbf{D}$ (we think of the two strategies as cooperation and defection, but this interpretation is irrelevant for our results). The choices of the two players determine their payoffs. Payoffs are assumed to be symmetric, and given by the (normalized) payoff matrix
[…]
We refer to this matrix as the stage game. As usual, player 1 chooses a row of this matrix, player 2 chooses a column, and the entries represent the payoffs of player 1. Depending on the signs of $A$ and $B$, we recover four major game classes: prisoner’s dilemma (A<0A\!<\!0, B>0B\!>\!0), stag-hunt game (A>0A\!>\!0, B>0B\!>\!0), snowdrift game (A<0A\!<\!0, B<0B\!<\!0), and harmony game (A>0A\!>\!0, B<0)B\!<\!0).
To describe how players adopt more profitable strategies over time, we use introspection dynamics [15, 16]. This process assumes that the two individuals interact in the same stage game over many time steps. Each time step, a randomly chosen player is given a chance to revise its strategy. That player then compares its current payoff $\pi$ with the payoff $\pi^{\prime}$ the player could have obtained by choosing the other strategy (keeping the co-player’s strategy fixed). As illustrated in Fig. 1A, the player switches to the other strategy with a probability that depends on the payoff difference [9, 55, 59],
@@
The parameter $\beta_{i}\!\geq\!0$ represents player $i$’s payoff sensitivity. For small $\beta_{i}$, this function approaches $1/2$ irrespective of the payoff difference (Fig. 1B left). In this limit, players decide whether to switch strategies based on a coin toss. For large $\beta_{i}$, the function approaches a step function (Fig. 1B right). Here, the player always switches if the alternative strategy yields the better payoff. We iterate this basic updating process for many time steps. As a result, players repeatedly get a chance to switch their strategies (Fig. 1C, bottom). Based on this process, we record how frequently we observe each of the four possible game outcomes over time. In the right panel of Fig. 1C, we illustrate these overall frequencies using a black-and-white gradient. A darker shading indicates game outcomes that are observed more often.
Introspection dynamics is particularly convenient because the expected frequencies of the four game outcomes can be computed analytically for any values of $\beta$ – contrary to other stochastic dynamics such as pairwise imitation [59] or experienced-weighted attraction learning models [13]. To compute the expected frequencies of each game outcome, we represent the learning dynamics as a Markov chain (see Supporting Information for details). The Markov chain has four states, corresponding to the four possible outcomes in each time step, CC, CD, DC, DD (the first letter refers to the first player’s strategy, and the second letter to the second’s). For finite $\beta_{1},\beta_{2}$, this Markov chain has a unique stationary distribution $\mathbf{u}(A,B,\beta_{1},\beta_{2})\penalty 10000\ :=\penalty 10000\ (u_{\textbf{CC}},u_{\textbf{CD}},u_{\textbf{DC}},u_{\textbf{DD}})$. This distribution captures how often we observe each game outcome on average, over the course of many rounds. Based on the stationary distribution, it is straightforward to compute player 1’s average payoff. To this end, we multiply the four frequencies $u_{ij}$ with the respective stage game payoffs,
@@
The respective results become particularly simple when both players have the same payoff sensitivity, $\beta_{1}\!=\!\beta_{2}$. In that case, one can show that the stationary distribution simplifies to
@@
For weak selection ($\beta\!\rightarrow\!0$), this formula implies that all game outcomes are observed equally often, $\mathbf{u}=(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})$. For strong selection ($\beta\!\rightarrow\!\infty$), the stationary distribution depends on the game parameters $A$ and $B$. For example, for the prisoner’s dilemma, (4) implies that the two players defect in almost all rounds ($u_{\textbf{DD}}\!=\!1$); in the snowdrift game, they anti-coordinate ($u_{\textbf{CD}}\!=\!u_{\textbf{DC}}\!=\!\frac{1}{2}$); and in the stag-hunt game they choose the strategy that is risk-dominant $u_{\textbf{DD}}\!=\!1$ if A<BA\!<\!B, $u_{\textbf{CC}}\!=\!1$ if $A>B$, see Ref. 15].
@@
Figure 2: Introspection dynamics among players with different payoff sensitivities. To illustrate the impact of payoff sensitivity on the learning dynamics, we consider two stage games, a prisoner’s dilemma (A–C) and a snowdrift game (D–F). In each case, we assume payoff sensitivities may take two possible values. They can either be comparably low (bright avatar) or high (dark avatar). For all possible combinations of payoff sensitivities, we depict how often the respective players would obtain one of the four possible payoffs of the respective stage game. Based on this stationary distribution, we compute the players’ expected payoffs. For the prisoner’s dilemma, we observe that the higher payoff sensitivity dominates the lower payoff sensitivity (indicated by blue arrows in panel D). In contrast, for the snowdrift game, no payoff sensitivity value is dominant. Instead, each player prefers to have the opposite payoff sensitivity value of their opponent.
Using the same formalism, we can also analyze players with different payoff sensitivities. In the Supporting Information, we derive a general expression of the stationary distribution $\mathbf{u}(A,B,\beta_{1},\beta_{2})$ as a function of $\beta_{1}$ and $\beta_{2}$, which is now more complex. In the following, we use that expression to explore whether a more sensitive learner (one with a larger $\beta$) would gain a long-term payoff advantage in various stage games.
In Fig. 2A–D, we consider a prisoner’s dilemma. We assume the two players can either have a small or a comparably large payoff sensitivity, $\beta\!\in\!\{2,5\}$. For a symmetric situation in which both players use the same $\beta$, we find that players learn to defect more often when $\beta$ is large (Fig. 2A,B). This result is intuitive: the larger the payoff sensitivity, the better players are at identifying the individually optimal strategy of defecting. Once players are asymmetric, we find that it is the more sensitive learner who defects more often (Fig. 2C). As a result, this player ends up with a better average payoff.
In Fig. 2E–H, we consider a snowdrift game. For illustration, the possible sensitivities are now $\beta\!\in\!\{2,16\}$. In the symmetric case among equally sensitive players, a large $\beta$ again makes it more likely that players implement a Nash equilibrium (here, CD or DC). In the depicted example, they coordinate on a Nash equilibrium in $63\%$ of all rounds when both have $\beta\!=\!2$, but in $91\%$ of all rounds when both have $\beta\!=\!16$. Remarkably, however, once players differ in their sensitivities, it is the less sensitive (and hence more erratic) learner who obtains the larger payoff (Fig. 2G). The intuition is the following. When in their less preferred equilibrium DC, the less sensitive player is still relatively likely to deviate. For the given payoff values, the respective switching probability according to Eq. [2] is $1/(1+{\rm e}^{2\times 0.1})=0.45$. This probability is smaller than $1/2$ (because the switch reduces the player’s payoff from 0.1 to 0), but it remains substantial. After this switch towards DD, the more sensitive player is tempted to react by playing C. As a result, the two players transitioned from one Nash equilibrium (DC, which favors the more sensitive player) to another (CD, which favors the less sensitive player). Interestingly, the converse transition from CD to DC is less likely because, starting from CD, the less sensitive player now has more to lose by deviating (from 1.5 to 1.0).
@@
Based on the above results, we can construct a ‘super-game’ matrix (Fig. 2D,H). This matrix represents the interaction of two learners who may have different payoff sensitivities. The entries of the matrix depict the learners’ average payoffs over the course of the learning process (as derived in the previous panels). For the prisoner’s dilemma, the super-game indicates that each learner always prefers to have a higher payoff sensitivity (indicated by blue arrows). In the snowdrift game example, the super-game takes the form of an anti-coordination game: players prefer their payoff sensitivity to be high when their opponent’s is low, and vice versa.
@@
Overall, these two examples provide a proof of principle. They illustrate the non-trivial effects of asymmetric payoff sensitivities. In general, the structure of the super-game depends on both the chosen $\beta$ values and the exact game parameters A and B. In particular, for the snowdrift game depicted in Fig. 2, one can also construct super-games in which the lower sensitivity dominates the larger one (or the other way round). Thus, these observations call for a more systematic analysis, considering more games and a larger set of possible sensitivities. We provide such an analysis in Fig. S1 of the Supporting Information. In addition, the above results raise the question of whether the players’ sensitivities would necessarily move towards ever-increasing values if it were an evolving trait. We explore that issue in the following.
Figure 3: Adaptive dynamics of payoff sensitivity for three different social dilemmas. We explore the adaptive dynamics for three different social dilemmas: the prisoner’s dilemma, the snowdrift game, and the stag-hunt game. For each case, we depict a realization of an individual-based simulation (upper panels) and a pairwise invasibility plot (lower panels). Each simulation is initialized at a monomorphic population with a payoff sensitivity of $\beta\!=\!0$. Over time, mutations introduce variation in the players’ payoff sensitivity. Individuals obtain payoffs by randomly interacting with other population members and learning strategies through introspection dynamics. They reproduce according to their payoff. For details, see Material and Methods. Pairwise invasibility plots illustrate the dynamics of monomorphic resident populations. They display (in color) which rare mutant traits ($y$-axis) have a positive invasion fitness $s_{x}(y)$, given the current resident ($x$-axis). Colored regions above the diagonal indicate that mutants with a higher trait value than the resident are favored to invade. Colored regions below the diagonal indicate evolution towards smaller trait values. Dashed lines indicate the position of singular trait values. A,B, For the depicted prisoner’s dilemma, we observe a dynamics of ever-increasing payoff sensitivities. C,D, In the snowdrift game, the evolving $\beta$ values converge to a finite value. E,F, In the stag-hunt game, we observe evolutionary branching. After this occurs, there are two subpopulations. Members of the first subpopulation have low $\beta$ values, and hence they choose strategies mostly at random. Members of the second subpopulation exhibit high $\beta$ values; they tend to best respond to their respective opponent.